A project costs $300 and has cash flows of $75 for each of the first three years and S50 in each of the projects next three please answe all
Answer #1

Solution4: Part c is correct- Around 4.5 years

YEAR CASH FLOW NET INVESTED CASH
0 -300
1 75 -225
2 75 -150
3 75 -75
4 50 -25
5 50 25
6 50
Pay back period No. of years required by the cash flow to recover initial cash outlay investment
Upto Year 4 :75+75+75+50=275 Recovered
After 4th year there is more cash inflow (50) to recover (-25) outstanding cash outlay
Pay back period =4+25/50=4.5

Solution 5: Part B: zero NPV

YEAR CASH FLOW

Present value of cash flows@i=5.84%

Formula:PV= FV/(1+R)^n

0 -2500
1 400 =400/1.0584=377.9289
2 400

=400/(1.0584)^2=357.0757

3 400

=400/(1.0584)^3=337.3731

4 400

=400/(1.0584)^4=318.7577

5 400 =400/(1.0584)^5=301.1694
6 400

=400/(1.0584)^6=284.5516

7 400

=400/(1.0584)^7=268.8507

8 400

=400/(1.0584)^8=254.0162

total present value of cash inflow 2499.723
total present value of cash inflow
2500 approx.
Present value of cash inflow=sum of the present value of each year cash inflow discounted at 5.84%
Present value of cash outflow=2500( Given)
Total Net present value=present value of cash inflow - present value of cash outflow =2500-2500=0